If sec theta + tan theta = x then sec theta equals to

Grade:11

1 Answers

sumit kumar pathak
6 years ago
sec theta + tan theta =xSo 1/cos theta + sin theta/cos theta =xSo cos theta =( 1+sin theta) /xHence sec theta = 1/cos theta =x/(1+sin theta)